OpenStudy (anonymous):

I need someone to check my work (algebra 2)

OpenStudy (anonymous):

can you paste the problems?

OpenStudy (anonymous):

Yes Hold on

OpenStudy (anonymous):

Let f(x) = -3x + 2, g(x) = x/5, h(x) = -2x^2 + 9, and j(x) = 5-x

OpenStudy (anonymous):